### v1

# Adaptive Data-Knowledge Alignment in Genetic Perturbation Prediction

Abstract
The transcriptional response to genetic perturbation reveals fundamental insights into complex cellular systems. While current approaches have made progress in predicting genetic perturbation responses, they provide limited biological understanding and cannot systematically refine existing knowledge. Overcoming these limitations requires an end-to-end integration of data-driven learning and existing knowledge. However, this integration is challenging due to inconsistencies between data and knowledge bases, such as noise, misannotation, and incompleteness. To address this challenge, we propose ALIGNED (Adaptive aLignment for Inconsistent Genetic kNowledgE and Data), a neuro-symbolic framework based on the Abductive Learning (ABL) paradigm. This end-to-end framework aligns neural and symbolic components and performs systematic knowledge refinement. We introduce a balanced consistency metric to evaluate the predictions’ consistency against both data and knowledge. Our results show that ALIGNED outperforms state-of-the-art methods by achieving the highest balanced consistency, while also re-discovering biologically meaningful knowledge. Our work advances beyond existing methods to enable both the transparency and the evolution of mechanistic biological understanding.
1 Introduction
Understanding how genetic perturbation affects transcriptional regulation is essential for deciphering complex biological systems, with profound implications for drug discovery and precision medicine (Badia-i Mompel et al., 2023; Gavriilidis et al., 2024; Ahlmann-Eltze et al., 2025). While advances in experimental technology now allow systematic interrogation of gene regulatory landscapes at an unprecedented scale (Norman et al., 2019; Replogle et al., 2022), existing datasets remain insufficient for building predictive models that can elucidate the full complexity of a cellular system (Peidli et al., 2024). This raises a critical question of how to design predictive frameworks that not only achieve high accuracy but also yield deeper biological understanding from these experimental capabilities.
Two complementary approaches have emerged, either by leveraging latent representations trained on extensive cell data (Lotfollahi et al., 2023; Theodoris et al., 2023; Cui et al., 2024; Hao et al., 2024) or incorporating prior biological knowledge for inductive biases (Roohani et al., 2024; Wang et al., 2024; Littman et al., 2025; Wenkel et al., 2025). Yet, both approaches provide limited insights into the biological mechanisms underlying their predictions. Data-driven models operate as black boxes, making it difficult to understand which regulatory relationships drive specific predictions (Bendidi et al., 2024). While hybrid methods incorporate prior biological knowledge, they treat this knowledge as static constraints rather than interpretable and updatable representations of biological understanding. Importantly, current approaches provide no end-to-end solution to identify and resolve divergences between data-driven learning and existing knowledge, which limits opportunities for continual refinement of biological understanding. (Gavriilidis et al., 2024; Kedzierska et al., 2025).
Overcoming these limitations requires explicitly integrating data-driven learning with established knowledge. However, a key challenge is the pervasive inconsistencies between experimental data and curated knowledge (Lu et al., 2024) due to imperfections in both information sources. Perturbation datasets exhibit multiple sources of noise (Liu et al., 2025; Rohatgi et al., 2024), experimental measurement biases (Kim et al., 2015; Peidli et al., 2024) and weak post-perturbation signals (Nadig et al., 2025; Aguirre et al., 2025). Meanwhile, transcriptional regulatory knowledge bases curated by experts often suffer from outdated information (Khatri et al., 2012), limited coverage (Saint-André, 2021) and biases towards better-studied pathways (Chevalley et al., 2025).
16.6%67.7%15.7%16.6%83.4%(a) OmniPath KB vs Norman15.3%70.7%14.0%15.3%84.7%(b) GO KB vs Norman8.2%42.4%49.4%8.2%91.8%(c) EcoCyc KB vs Precise1KConsistent (Inner)Inconsistent (Inner)Missing in KB (Outer)Data-KB Conflict (Outer)
Figure 1: Inconsistency between gene regulatory knowledge bases (KBs) and data-derived perturbation-responses correlations. We examined OmniPath (Türei et al., 2016), Gene Ontology (GO) (Ashburner et al., 2000) and EcoCyc (Moore et al., 2024) knowledge bases, human (Norman et al., 2019) and bacterial (Precise1k, Lamoureux et al., 2023) datasets.
To illustrate this challenge, we analyzed popular knowledge bases and benchmark datasets (Figure 1), finding that 42-71% of data-derived regulatory relationships are missing across curated knowledge bases, while a minimum of 14% directly conflict with existing annotations. Naive integration of inconsistent sources risks bidirectional error propagation (Lu et al., 2024) that can corrupt both data-driven learning and knowledge refinement. This inconsistency prevents models from effectively leveraging prior biological knowledge in predictions (Ahlmann-Eltze et al., 2025) and compromises their ability to produce biologically meaningful regulatory relationships from learned representations.
To address this challenge, the Abductive Learning (ABL) paradigm (Zhou, 2019; Huang et al., 2023) offers a foundation for integrating data-driven learning with symbolic knowledge refinement through consistency optimization. Based on this approach, we propose ALIGNED (Adaptive aLignment for Inconsistent Genetic kNowledgE and Data), an end-to-end framework that enables neuro-symbolic alignment and knowledge refinement in genetic perturbation prediction. ALIGNED advances beyond existing predictive methods to enhance transparency about the underlying biological mechanisms and enable continual evolution of understanding from large-scale perturbation datasets.
Our main contributions are:
  • Balanced Consistency Metric. We design a balanced evaluation metric that assesses predictions against both experimental data and curated knowledge. This addresses the limitation that standard metrics evaluate only predictive accuracy without considering consistency with biological knowledge (Bendidi et al., 2024).
  • Adaptive Neuro-Symbolic Alignment. We align neural and symbolic predictions from inconsistent information sources by adaptively weighting neural and symbolic components with a gradient-free optimization mechanism.
  • Knowledge Refinement. We enable systematic update of regulatory interactions by introducing a gradient-based optimization approach over a symbolic representation of the GRNs.
  • Results. ALIGNED outperforms existing methods in balanced consistency with both data and knowledge. In addition, ALIGNED’s knowledge refinement can re-discover cross-referenced regulatory relationships. Our results demonstrate effective translation from prediction to enhanced mechanistic interpretation.
2 Preliminaries
2.1 Problem Setting
We formalize the prediction of genome-scale response to genetic perturbation as a ternary classification problem. The goal is to learn a function $f:\{-1,0,1\}^{n}\rightarrow\{-1,0,1\}^{n}$, where $n$ is the total number of genes. The input values 1-1, 0, and $1$ represent negative perturbation (deletion or knockout), no perturbation, and positive perturbation (overexpression), respectively. The output values indicate decreased expression, no significant change, or increased expression for each gene.
We denote the labelled dataset by $D_{l}=\langle{\bm{X}}_{l},{\bm{Y}}_{l}\rangle$, where ${\bm{X}}_{l}$ contains the perturbed gene inputs and ${\bm{Y}}_{l}$ contains the corresponding perturbation responses obtained from transcriptome sequencing experiments. An unlabelled dataset ${\bm{X}}_{u}$ is also used for training in abductive learning, which contains only the perturbation input.
2.2 Symbolic reasoning over gene regulatory networks
We explain symbolic reasoning methods that allow us to predict perturbation responses. We focus on gene regulatory networks (GRNs) as our knowledge bases, which contain activation (+) and inhibition (-) interaction relations between genes. We utilize symbolic reasoning via Boolean matrices (Ioannidis & Wong, 1991; Ai, 2025). Direct activation and inhibition interactions are compiled as $n\times n$ adjacency matrices $\langle{\bm{R}}_{+}^{(0)},{\bm{R}}_{-}^{(0)}\rangle$:
${\bm{R}}_{+}^{(i)}={\bm{R}}_{+}^{(0)}\cdot{\bm{R}}_{+}^{(i-1)}+{\bm{R}}_{-}^{(0)}\cdot{\bm{R}}_{-}^{(i-1)}$
${\bm{R}}_{-}^{(i)}={\bm{R}}_{+}^{(0)}\cdot{\bm{R}}_{-}^{(i-1)}+{\bm{R}}_{-}^{(0)}\cdot{\bm{R}}_{+}^{(i-1)}$ (1)
We approximate the fixpoint of $\langle{\bm{R}}_{+}^{(\infty)},{\bm{R}}_{-}^{(\infty)}\rangle$ by interleaving the computations with respect to a partial ordering on the matrices ${\bm{R}}_{+}^{(k)},{\bm{R}}_{-}^{(k)}$ for a finite $k$ (Tarski, 1955). The obtained knowledge base $\mathcal{KB}=\langle{\bm{R}}_{+}^{(k)},{\bm{R}}_{-}^{(k)}\rangle$ represents indirect regulations via pathways up to a maximum length of $k$ interactions.
Given an input perturbation $\bm{x}$}, we infer its effect on a genome scale by performing a deductive query in the knowledge base $\mathcal{KB}$. The matrix operations $\delta_{\mathcal{KB}}({\bm{x}})=({\bm{R}}_{+}^{(k)}-{\bm{R}}_{-}^{(k)})^{\top}{\bm{x}}$ allow us to perform this query with high computational efficiency. Based on this approach, we define a measurement for the data-knowledge inconsistency illustrated in Figure 1:
$\mathrm{Inc}(D_{l},\mathcal{KB})=\sum\limits_{{\bm{x}},{\bm{y}}\in D_{l}}\|\delta_{\mathcal{KB}}({\bm{x}})-{\bm{y}}\|_{0}$ (2)
where $D_{l}=\langle{\bm{X}}_{l},{\bm{Y}}_{l}\rangle$ is a labelled dataset. Our approach differs from the Known Relationships Retrieval metric (Celik et al., 2024; Bendidi et al., 2024) in that the deductive queries respect the global GRN structure and preserve the transitivity of genetic interactions.
2.3 Abductive Learning
Our framework explicitly integrates the neural and symbolic components and handles data-knowledge inconsistencies based on the Abductive learning (ABL) paradigm (Zhou, 2019). ABL is a neuro-symbolic approach that aims to learn a function $f$ and align its predictions with the knowledge base $\mathcal{KB}$ via consistency optimization.
A general ABL training pipeline takes a neural model $f$ pretrained on labelled data $\langle{\bm{X}}_{l},{\bm{Y}}_{l}\rangle$ as initialization. From the unlabelled dataset ${\bm{X}}_{u}$, $f$ makes neural predictions $\hat{{\bm{y}}}=f({\bm{x}}_{u})$, which may be inconsistent with $\mathcal{KB}$. Consistency optimization is then performed, with revising $\hat{{\bm{y}}}$ to $\bar{{\bm{y}}}$ and updating $f$ on the revised dataset $\langle{\bm{X}}_{u},\bar{{\bm{Y}}}\rangle$. This process can be executed iteratively until convergence or reaching an iteration limit $T$. Formally,
$f:{\bm{X}}_{l}\rightarrow{\bm{Y}}_{l}$
$\mathrm{s.t.}\quad$ $\forall{\bm{x}}\in{\bm{X}}_{l}\cup{\bm{X}}_{u}:\ \mathcal{KB}\models\langle{\bm{x}},f({\bm{x}})\rangle,\ \mathrm{or}$
$f({\bm{x}})=\delta_{\mathcal{KB}}({\bm{x}}),\mathcal{KB}\models\langle{\bm{x}},\delta_{\mathcal{KB}}({\bm{x}})\rangle$
where $\delta({\bm{x}},\mathcal{KB})$ is the symbolic prediction on $\bm{x}$} by $\mathcal{KB}$ and $\models$” denotes logical entailment.
3 The ALIGNED Method
Figure 2: The ALIGNED (Adaptive aLignment of Inconsistent Genetic kNowledgE and Data) framework. ALIGNED contains a neural component (blue), a symbolic component (green) and an adaptor (purple).
We first introduce a consistency metric for evaluating how well predictions align with both the test data and the knowledge base. We then present ALIGNED (Adaptive aLignment for Inconsistent Genetic kNowledgE and Data), a framework which adaptively integrates reliable information from both sources to predict genetic perturbation responses.
3.1 The Balanced Consistency Metric
To evaluate the consistency of a prediction against both the test data and the knowledge base, we define a balanced consistency metric $F_{1\ \mathrm{balance}}$, which considers the $F_{1}$ scores from both the test dataset and the knowledge base. $F_{1\ \mathrm{balance}}$ includes a coefficient $\gamma>1$ to balance the two $F_{1}$ scores and penalize when either score being too low:
$F_{1\ \mathrm{balance}}(f({\bm{x}}),{\bm{x}},{\bm{y}},\mathcal{KB})=$ $\big(\frac{1}{2}F_{1}({\bm{y}},f({\bm{x}}))^{-\gamma}+\frac{1}{2}F_{1}(\delta_{\mathcal{KB}}({\bm{x}}),f({\bm{x}}))^{-\gamma}\big)^{-1/\gamma}$ (3)
3.2 ALIGNED Framework Overview
The ALIGNED framework (Figure 2) integrates three components to balance data-knowledge inconsistencies through iterative refinement. The neural component $f_{y}$ is a neural network which predicts perturbation responses from input data, while the symbolic component $\mathcal{KB}$ performs symbolic reasoning over gene regulatory networks encoded as matrices (computed via Equation 2.2). The adaptor $f_{a}$ learns to combine neural and symbolic predictions based on their relative reliability for each prediction.
Training proceeds using both labelled and unlabelled data. We initialize components $f_{y}$ and $f_{a}$ by training them jointly on the labelled dataset. For each unlabelled input, the framework produces a neural and a symbolic prediction. In adaptive alignment, since these predictions may be inconsistent, the adaptor is trained to produce a binary indicator vector that selects which predictive source to trust for each output dimension. This creates an integrated neuro-symbolic prediction that combines results from both predictive sources. The framework then performs multiple iterations of alignment and bidirectional updates to neural and symbolic components. Using the neural-symbolic predictions, we re-train the neural component and perform knowledge refinement to the symbolic component.
3.3 Adaptive Neuro-Symbolic Alignment with Gradient-Free Optimization
In this section, we will introduce the alignment mechanism used by ALIGNED to adaptively integrate neural and symbolic predictions. We denote a binary alignment indicator vector ${\bm{a}}=f_{a}({\bm{x}})$ and the neuro-symbolic prediction $\bar{{\bm{y}}}$. After the initialization and each round of bidirectional update, $\bar{{\bm{y}}}$ is produced from both neural prediction $\hat{{\bm{y}}}$ and symbolic prediction $\delta_{\mathcal{KB}}({\bm{x}})$, according to the indicator $\bm{a}$} such that neural prediction is used when ${a}_{i}=0$ and symbolic prediction when ${a}_{i}=1$.
Our definition of the training objectives for the adaptor can be divided into three parts. First, since information derived from experimental data and curated knowledge may be inconsistent, the adaptor considers how neural-symbolic predictions differ from the curated knowledge. We describe this using the inconsistency between $\bar{{\bm{y}}}$ and $\mathcal{KB}$ based on Equation 2:
$\mathrm{Inc}({\bm{a}},{\bm{x}},\hat{{\bm{y}}},\mathcal{KB})=\|\delta_{\mathcal{KB}}({\bm{x}})-\bar{{\bm{y}}}\|_{0}$
Second, we design a loss term to leverage as much information from labelled training data as possible to reduce the predictions’ inconsistency with data. Therefore, we restrict the framework to only use knowledge-derived information when necessary. We defined this restriction with threshold $\theta$:
$L_{len}({\bm{a}})=\max\{\|{\bm{a}}\|_{0}-\theta,0\}$
Third, we take into account how well each gene is represented in knowledge. To measure this, we use a weight vector $\bm{w}$} as hyper-parameter, which contains the number of training data samples that are inconsistent with $\mathcal{KB}$ (computed by Equation 2), and the number of annotations from Gene Ontology (Ashburner et al., 2000). This allows us to reward $\bm{a}$} by maximizing the usage of symbolic prediction when a gene is represented well in $\mathcal{KB}$, otherwise use neural prediction. We define a loss with regard to $\bm{w}$}:
$L_{weight}({\bm{a}})={\bm{w}}^{\top}(\bm{1}-{\bm{a}})+(\bm{1}-{\bm{w}})^{\top}{\bm{a}}$
where higher values of ${w}_{i}$ indicate that gene $i$ is well-represented in the knowledge base and more consistent with data, suggesting the symbolic prediction should be preferred. Lower values indicate sparser knowledge or more data-knowledge conflicts, and so the neural prediction should be favored. We combine the above three parts in the adaptor’s objective:
[…]
where hyper-parameter $C_{w}$, $C_{l}$ are trade-off coefficients. Minimizing $L_{a}$ includes querying the symbolic $\mathcal{KB}$, which has a discrete structure. This creates a combinatorial optimization problem, so a gradient-free optimization method is necessary. We train $f_{a}$ with the REINFORCE algorithm (Williams, 1992; Hu et al., 2025) and initialize its sampling distribution based on $\bm{w}$} to reduce sampling complexity. To exploit representations captured by the neural component $f_{y}$ from the experimental data, $f_{a}$ shares input $\bm{x}$} and embedding layers with $f_{y}$. We optimize $f_{y}$ and $f_{a}$ jointly with the following objective:
$\min\limits_{f_{y},f_{a}}\ \mathcal{L}=\,$ $\frac{1}{|D_{l}|}\sum\limits_{({\bm{x}},{\bm{y}})\in D_{l}}CE(f_{y}({\bm{x}}),{\bm{y}})$
$+$ $C\frac{1}{|D_{l}\cup D_{u}|}\sum\limits_{{\bm{x}}\in D_{l}\cup D_{u}}L_{a}({\bm{a}},{\bm{x}},\hat{{\bm{y}}})\log f_{a}({\bm{x}})$ (5)
[…]
3.4 Gradient-Based Knowledge Refinement with Sparse Regularization
To address missing and inaccurate interactions in $\mathcal{KB}$, we incorporate a knowledge refinement mechanism into the ALIGNED framework that leverages reliable information from neural and symbolic predictions. For computational efficiency on large-scale GRNs, we consider gradient-based optimization, and introduce an approximation function $\varepsilon(\cdot)$ for Boolean elements (Ravanbakhsh et al., 2016). This approximation enables gradient-based optimization compatibility of the non-differentiable Boolean matrix multiplication in Equation 2.2:
$\varepsilon_{t}({\bm{X}})_{i,j}=1-\exp(-t{X}_{i,j}),{X}_{i,j}\geq 0$
We introduce an inductive bias for minimal modifications to the GRN during refinement. This ensures the biological relationships and structure in the GRN are not distorted by noise in the data. We perform an $l_{1}$ sparse regularizied optimization to achieve this, fitting $\mathcal{KB}$ to neuro-symbolic predictions using proximal gradient descent (Tibshirani, 1996; Candes & Recht, 2012). The objective of knowledge refinement is defined as follows:
$\min\limits_{{\bm{P}}^{(0)}_{+},{\bm{P}}^{(0)}_{-}\in\mathbb{R}_{+}^{n\times n}}\ \mathcal{L}_{\mathrm{refine}}({\bm{P}}^{(0)}_{+},{\bm{P}}^{(0)}_{-},k)=$ $\sum\limits_{{\bm{x}},{\bm{y}}\in\langle{\bm{X}}_{u},\bar{{\bm{Y}}}\rangle}\|\varepsilon_{t_{k}}({\bm{P}}_{+}^{(k)}-{\bm{P}}_{-}^{(k)})^{\top}{\bm{x}}-{\bm{y}}\|_{2}^{2}$
$+\lambda\big(\|\varepsilon_{t_{0}}({\bm{P}}_{+}^{(0)})-{\bm{R}}_{+}^{(0)}\|_{1}+\|\varepsilon_{t_{0}}({\bm{P}}_{-}^{(0)})-{\bm{R}}_{-}^{(0)}\|_{1}\big)$ (6)
where ${\bm{R}}^{(0)}_{+}$ and ${\bm{R}}^{(0)}_{-}$ represent the initial GRN before refinement, real-valued non-negative matrices ${\bm{P}}_{+}^{(0)}$ and ${\bm{P}}_{-}^{(0)}$ are the refined GRN with direct regulatory interactions. To facilitate gradient passing, we use real-number matrix computation instead of Boolean matrix in Equation 2.2 and use ${\bm{P}}_{+}^{(0)}$ and ${\bm{P}}_{-}^{(0)}$ to compute indirect regulatory interactions ${\bm{P}}_{+}^{(k)}$ and ${\bm{P}}_{-}^{(k)}$. $\lambda$ denotes a regularization parameter, $t_{k},t_{0}$ denote coefficients of approximation, $\|\cdot\|_{1}$ denotes the element-wise matrix $l_{1}$ norm.
4 Experiments
We evaluate ALIGNED on multiple large-scale perturbation datasets for predicting genome-wide responses and assess the knowledge refinement mechanism in isolation. The experiments address the following research questions:
  • Q1
    Can ALIGNED achieve a higher balanced consistency than existing methods without damaging either data or knowledge consistency?
  • Q2
    Is the knowledge refinement mechanism capable of re-discovering biologically meaningful and well-structured regulatory knowledge?
  • Q3
    Does the framework leverage knowledge to improve prediction on unseen data, particularly under limited data availability?
4.1 Perturbation Prediction on Benchmark Datasets (Q1)
We focused on multiple large-scale perturbation datasets that are widely adopted for this prediction task: 1) Norman et al. (2019) for human K562 cells, including gene expression profiles under single and double perturbations across 102 genes (128 double and 102 single perturbations) with 89,357 samples; 2) Dixit et al. (2016) for mouse BDMC cells, containing 19 single gene perturbations with 43,401 samples; and 3) Adamson et al. (2016) for human K562 cells, containing 82 single gene perturbations with 65,899 samples. Our knowledge base $\mathcal{KB}$ integrates the Omnipath GRN (Türei et al., 2016) and the GO-based gene interaction graph from Roohani et al. (2024), covering 3,949 genes for the Norman et al. (2019) dataset and 2,958 genes for the Dixit et al. (2016) dataset. To evaluate methods on unseen perturbations, we split the test set of Norman et al. (2019) dataset to include 19 unseen single-gene perturbations and 18 unseen double-gene perturbations. The Dixit et al. (2016) and Adamson et al. (2016) datasets were split randomly.
We evaluated ALIGNED variants built with an MLP or a $\mathcal{KB}$-embedded GNN as the neural component. Performance of ALIGNED was compared in Figure 3 with state-of-the-art methods including: 1) GEARS, a GNN-based data-knowledge hybrid model (Roohani et al., 2024); 2) foundation models scGPT (Cui et al., 2024) and scFoundation Hao et al. (2024); 3) a linear additive perturbation model incorporating regulatory knowledge (Ahlmann-Eltze et al., 2025). To ensure that all methods are measured on the same knowledge base, ALIGNED does not perform knowledge refinement during this comparison.
To assess the contribution of each framework component, we conducted an ablation study comparing the complete ALIGNED framework against its neural component baseline (trained only on labelled data) in Figure 4. We tracked performance through two complete iterations of ALIGNED, with each iteration consisting of adaptive neuro-symbolic alignment followed by knowledge refinement (denoted as Align/Refine 1, 2).
For each method, we evaluated data consistency $F_{1}(\bar{{\bm{Y}}},{\bm{Y}}_{test})$ which measures performance of the predictions, knowledge consistency $F_{1}(\bar{{\bm{Y}}},\delta_{\mathcal{KB}}({\bm{X}}))$, and the balanced consistency metric $F_{1\ \mathrm{balance}}$ as defined in Equation 3.
Balanced Cons.Data Cons.Knowledge Cons.0.30.350.40.450.50.55(a) Norman et al. DatasetBalanced Cons.Data Cons.Knowledge Cons.0.10.20.30.40.5(b) Dixit et al. DatasetBalanced Cons.Data Cons.Knowledge Cons.0.20.30.40.50.6(c) Adamson et al. DatasetALIGNED (GNN)ALIGNED (MLP)LinearGEARSscGPTscFoundation
Figure 3: Balanced, data and knowledge consistency across methods.
[…]
Figure 4: Performance of the complete ALIGNED framework built with GNN.
Observation 1. In Figure 3, ALIGNED achieved significantly higher knowledge consistency than other methods, with slightly higher data consistency. It consequently outperformed existing methods in balanced consistency. This shows ALIGNED’s ability to make a better trade-off between inconsistent data and knowledge, enabling the framework to provide mechanistic understandings for black-box neural predictions.
Observation 2. In Figure 4, after one round of alignment and refinement, ALIGNED improved knowledge consistency significantly while keeping comparable data consistency. This further demonstrates that ALIGNED had learned an effective adaptor function to trade off data- and knowledge-derived information.
4.2 Knowledge Refinement of Gene Regulatory Networks (Q2)
In this section, we aim to answer whether ALIGNED can re-discover biologically meaningful and well-structured knowledge. We tested ALIGNED’s knowledge refinement in isolation and evaluated the refined GRN interactions in three aspects: accuracy (Figure 5a), topology (Figure 5b) and pathway enrichment (Figure 5c).
We used the accuracy of interactions to test if ALIGNED’s knowledge refinement can re-discover underlying regulations from synthetic data generated from OmniPath GRN (Türei et al., 2016). For topology, we evaluated the method’s ability to produce well-structured GRNs, in terms of: 1) network modularity, for clustering quality of functional modules (Alon, 2007) and 2) degree assortativity, for regulatory hub structures (Segal et al., 2003). In addition, we examined the method’s ability to re-discover biologically meaningful interactions by cross-referencing external pathway databases. We compared overlaps with refined pathways using a gene set recovery algorithm (Huang et al., 2018) to obtain pathway enrichment scores.
The original OmniPath GRN (Türei et al., 2016) contains 2,958 genes and 113,056 regulatory interactions. We corrupted the original GRN by randomly adding and removing equal numbers of interactions at different noise levels, ranging from 5% to 90%, to simulate varying degrees of knowledge base errors. The experiment aimed to recover the original GRN from its synthetic data, using our knowledge refinement method initialized with the noisy GRN.
Our method was compared with a baseline using non-sparse (Frobenius norm) regularization. Existing approaches, such as GRN inference, treat the knowledge base as static instead of performing incremental refinement, and therefore are not suitable for the comparison.
The accuracy was measured in $F_{1}$ score on both direct and indirect interactions (defined as Equation 2.2) of refined GRNs, assuming the original GRN as ground-truth. Network modularity and assortativity were measured on direct interactions of the GRN, with higher modularity scores for better clustering quality, and assortativity is usually negative in GRNs with well-structured regulatory hubs. To show the method’s ability in re-discovering biologically meaningful interactions, we took 302 pathways from the KEGG pathway database (Kanehisa et al., 2025) as a cross-reference for gene set recovery, and measured the difference of recovery scores between reconstructed and original GRN for each pathway.
[…]
Figure 5: GRN knowledge refinement performance with ALIGNED.
Observation 1. In Figure 5a, the accuracy of refined interactions by ALIGNED remained high ($F_{1}>0.7$) even with up to 40% noise. This shows that underlying regulatory knowledge in synthetic data can be captured by ALIGNED.
Observation 2. In Figure 5b, up to 20% noise, the topological measurements of the refined interactions are similar to those from the original GRN. This demonstrates the ability of ALIGNED in producing well-structured refined GRNs.
Observation 3. In Figure 5c, there are no significant differences of enrichment scores between the original and refined GRNs in most pathways. This indicates that ALIGEND can re-discover biologically meaningful knowledge annotated in cross-reference databases.
4.3 Perturbation Prediction on Bacterial Genome (Q3)
Setting, Dataset, and Knowledge Base. We evaluate our method on the Escherichia coli (E. coli) K-12 MG1655 strain using a combined dataset that includes 70 knockout perturbations by Lamoureux et al. (2023) (comprising 4 triple, 7 double, and 59 single perturbations, totaling 433 samples); and 7 data series with 16 single overexpression perturbations (73 samples) from the NCBI sequence read archive (Sayers et al., 2025). The knowledge base is constructed from the EcoCyc GRN (Moore et al., 2024), covering 315 regulator genes and 3,004 regulated genes. To evaluate generalization on unseen instances, we split the test set of unseen perturbations including 4 single overexpressions, 5 double knockouts, and 2 triple knockouts.
Similar to Section 4.1, we conducted ablation studies comparing ALIGNED against baseline models trained only on labelled data. We tracked performance through two complete iterations of ALIGNED to assess the cumulative effect of each framework component.
Observation 1. In Table 1, performance of prediction, i.e. data consistency, was significantly improved on unseen perturbations, simultaneously improving knowledge and balanced consistency. This indicates ALIGNED’s ability of effectively leveraging knowledge-derived information under limited data availability.
Table 1: Performance of ALIGNED on E.coli genome.
Model ABL Stage Data Cons. Knowledge Cons. Balanced Cons.
MLP Baseline $0.3312_{\pm 0.0004}$ $0.3371_{\pm 0.0009}$ $0.3341_{\pm 0.0006}$
GNN Baseline $0.3773_{\pm 0.0026}$ $0.3605_{\pm 0.0024}$ $0.3689_{\pm 0.0012}$
ALIGNED
Align 1 $0.3872_{\pm 0.0009}$ $0.3708_{\pm 0.0006}$ $0.3784_{\pm 0.0002}$
Refine 1 $0.3872_{\pm 0.0009}$ $0.3836_{\pm 0.0067}$ $0.3851_{\pm 0.0033}$
Algin 2 $0.3812_{\pm 0.0029}$ $0.4025_{\pm 0.0020}$ $0.39404_{\pm 0.0018}$
Refine 2 $0.3812_{\pm 0.0029}$ $0.4045_{\pm 0.0038}$ $0.3912_{\pm 0.0013}$
ALIGNED
Algin 1 $0.3876_{\pm 0.0060}$ $0.3714_{\pm 0.0269}$ $0.3800_{\pm 0.0159}$
Refine 1 $0.3876_{\pm 0.0060}$ $0.4520_{\pm 0.0070}$ $0.4130_{\pm 0.0059}$
Align 2 $0.3878_{\pm 0.0064}$ $0.4668_{\pm 0.0235}$ $0.4124_{\pm 0.0042}$
Refine 2 $\mathbf{0.3878}_{\pm 0.0064}$ $\mathbf{0.5348}_{\pm 0.0069}$ $\mathbf{0.4288}_{\pm 0.0063}$
5 Related Work
Perturbation Response Prediction. Recent approaches fall into two categories: methods that utilize the compositional nature of genetic perturbation responses in learning latent representations (Lotfollahi et al., 2023; Cui et al., 2024; Hao et al., 2024), hybrid methods that leverage prior knowledge from biological networks (Roohani et al., 2024; Wenkel et al., 2025; Littman et al., 2025) or textual embeddings (Wang et al., 2024). In contrast, ALIGNED does not assume GRNs to be static and can systematically refine GRNs by adaptive learning from datasets and knowledge bases to leverage reliable information.
Neuro-Symbolic Learning. The Abductive Learning (ABL) framework (Zhou, 2019) integrates deep learning with symbolic constraints through consistency optimization. Extensions include $Meta_{abd}$ (Dai & Muggleton, 2021) for visual-symbolic reasoning, $ABL_{NC}$ for knowledge refinement (Huang et al., 2023), and $ABL_{refl}$ (Hu et al., 2025) for efficient neuro-symbolic integration using reinforcement learning mechanisms. Additionally, Cornelio et al. (2023) proposed a learnable trade-off mechanism between data and knowledge sources. We extend these approaches to biological systems where both experimental data and curated knowledge exhibit domain-specific noise and incompleteness.
6 Conclusion and Future Work
In this work, we introduced ALIGNED, a novel end-to-end framework that achieves balanced neuro-symbolic alignment and knowledge refinement for predicting genetic perturbation. Importantly, our work not only enhances transparency about the biological relationships behind predictions but also enables the evolution of biological knowledge from large-scale datasets, advancing beyond current black-box approaches.
While we acknowledge the limitations in our regulatory network modelling, alternative methods (Covert et al., 2004; Stoll et al., 2017; Abou-Jaoudé et al., 2016) face significant scalability issues. Future work could explore differentiable models (Faure et al., 2023) and refine them with experimental data. Furthermore, ALIGNED can be extended to different biological tasks by leveraging other prior knowledge, such as protein-protein interaction networks (Rodriguez-Mier et al., 2025) and metabolic networks (Faure et al., 2023).
Appendix A Usage of Large Language Models
We used Claude and ChatGPT mainly to polish the language after all intellectual content has been drafted, along with other language editing tools such as Grammarly.
Appendix B Details of the Knowledge Base
B.1 Formal Definition
With regard to positivity of edges, the GRN can be represented as a datalog program (Minker, 1988) with two predicates, $\mathit{regulates_{+}}/2$ and $\mathit{regulates_{-}}/2$, where genes are constant names. We use the transitive closure of $\mathit{regulates_{+}}$ and $\mathit{regulates_{-}}$ as the knowledge base in the reasoning component of our framework, denoted as $r_{+}/2$ and $r_{-}/2$. The transitive closure can be evaluated as a bilinear recursive program (Ioannidis & Wong, 1991):
$r_{+}\leftarrow(\mathit{regulates_{+}}(a,c)\wedge r_{+}(c,b))\vee(\mathit{regulates_{-}}(a,c)\wedge r_{-}(c,b))$
$r_{-}\leftarrow(\mathit{regulates_{+}}(a,c)\wedge r_{-}(c,b))\vee(\mathit{regulates_{-}}(a,c)\wedge r_{+}(c,b))$
$r_{+}(a,b)\leftarrow\mathit{regulates_{+}}(a,b)$
$r_{-}(a,b)\leftarrow\mathit{regulates_{-}}(a,b)$ (7)
The transitive closure represents all regulatory pathways, accounting for the indirect effects of positive and negative regulation. The recursive program states that all direct positive (negative) regulations are included in the positive (negative) transitive closure, while an indirect pathway containing an even number of negative regulations contributes to the positive closure, and an odd number of negative regulations contributes to the negative closure. The datalog program can then be compiled as a recursive Boolean matrix multiplication (Equation 2.2), where matrices of positive (negative) direct regulations ${\bm{R}}_{+}^{(0)},{\bm{R}}_{-}^{(0)}\in\{0,1\}^{n\times n}$ are compiled from $\mathit{regulates}_{+},\mathit{regulates}_{-}$:
$({R}^{(0)}_{+})_{ij}=\begin{cases}1,\quad\mathit{regulates_{+}}(i,j)\\ 0,\quad\mathrm{otherwise}\end{cases},\ ({R}^{(0)}_{-})_{ij}=\begin{cases}1,\quad\mathit{regulates_{-}}(i,j)\\ 0,\quad\mathrm{otherwise}\end{cases}$
and indirect regulation matrices ${\bm{R}}_{+}^{(k)},{\bm{R}}_{-}^{(k)}$ are computed from Equation 2.2.
B.2 Demonstration
For a 5 nodes example GRN in Figure 6, a simple demonstration of the approximative fixpoint of regulatory interactions (Equation 2.2) is shown as Figure 7.
In this example, ${\bm{R}}_{+}^{(0)},{\bm{R}}_{-}^{(0)}$ are compiled as:
${\bm{R}}_{+}^{(0)}=\begin{bmatrix}1&0&1&0&1\\ 0&1&0&0&1\\ 0&0&1&1&0\\ 0&0&0&1&0\\ 0&0&0&0&1\end{bmatrix},{\bm{R}}_{-}^{(0)}=\begin{bmatrix}0&1&0&0&0\\ 0&0&0&0&0\\ 0&0&0&0&0\\ 1&0&0&0&1\\ 0&0&0&0&0\\ \end{bmatrix}$,
and the symbolic prediction when $k=2$ is:
$\delta_{\langle{\bm{R}}_{+}^{(2)},{\bm{R}}_{-}^{(2)}\rangle}({\bm{x}})=\big(\begin{bmatrix}1&0&1&1&1\\ 0&1&0&0&1\\ 0&1&1&1&0\\ 0&1&0&1&1\\ 0&0&0&0&1\end{bmatrix}-\begin{bmatrix}1&1&0&0&1\\ 0&0&0&0&0\\ 1&0&1&0&1\\ 1&0&1&1&1\\ 0&0&0&0&0\\ \end{bmatrix}\big)^{\top}\begin{bmatrix}1\\ 0\\ 0\\ 0\\ 0\end{bmatrix}=\begin{bmatrix}0\\ -1\\ 1\\ 1\\ 0\end{bmatrix}$
And $\delta_{\langle{\bm{R}}_{+}^{(k)},{\bm{R}}_{-}^{(k)}\rangle}({\bm{x}})=\bm{0}$ for $k\leq 4$ in this example, due to the negative feedback loop.
B.3 Discussion
In actual experiments in Section 4.1, we introduced an assumption that up/down regulation of a node can be decided by its in-degree of positive and negative interactions, in order to capture more detailed regulatory behaviours. This is achieved by using an integer variant of Equation 2.2. In this setting, the value of node 5 in Figure 7 will be -1 when $k=2$, i.e. $\delta_{\langle{\bm{R}}_{+}^{(2)},{\bm{R}}_{-}^{(2)}\rangle}({\bm{x}})=[0,-1,1,1,-1]^{\top}$, and the network behaviour will be more complicated in complex networks. However, such modelling is still not enough to describe the biological reality, and future work could further explore the other differentiable modelling approaches of genome-scale GRN under the ALIGNED framework.
Figure 6: A 5-node GRN example.
Figure 7: A demonstration of approximative fixpoint.
Appendix C Framework Overview
Additional demonstrations for the ALIGNED framework, including an overview figure Figure 8 and pseudo-code Algorithm 1, are included here.
Figure 8: The ALIGNED framework. The nueral component is marked in blue, and the symbolic component ingreen.
input : Labelled dataset $D_{l}=({\bm{X}}_{l},{\bm{Y}}_{l})$; Unlabelled dataset $D_{u}={\bm{X}}_{u}$;
[…]
output : Trained neural model $(f_{y},f_{a})$;
Updated knowledge base $\mathcal{KB}$
1 $\mathcal{KB}\leftarrow\langle{\bm{R}}_{+}^{(k)},{\bm{R}}_{-}^{(k)}\rangle$;
2 $(f_{y},f_{a})\leftarrow\mathrm{train}({\bm{X}}_{l},{\bm{Y}}_{l},\mathcal{KB})$;
3 for $1\leq t\leq T$ do
4 $\hat{{\bm{Y}}}\leftarrow f_{y}({\bm{X}}_{u})$;
5 ${\bm{A}}\leftarrow f_{a}({\bm{X}}_{u})$;
6 $\delta_{\mathcal{KB}}\leftarrow{\bm{X}}_{u}({\bm{R}}_{+}^{(k)}-{\bm{R}}_{-}^{(k)})$;
7 $\bar{{\bm{Y}}}\leftarrow\begin{cases}\bar{{Y}}_{ij}=\hat{{Y}}_{ij},&{A}_{ij}=0\\ \bar{{Y}}_{ij}=\delta_{\mathcal{KB}}({\bm{X}}_{u})_{ij},&{A}_{ij}=1\end{cases}$;
8 $(f_{y},f_{a})\leftarrow\mathrm{train}({\bm{X}}_{u},\bar{{\bm{Y}}}_{u})$;
9 $\mathcal{KB}\leftarrow\mathrm{refine}({\bm{X}}_{u},\bar{{\bm{Y}}}_{u},\mathcal{KB})$
10 end for
Algorithm 1 ALIGNED Framework
Appendix D Experiment Details
D.1 Scalability
All experiments Figure 3 were conducted on a Slurm-managed Linux cluster equipped with Intel Xeon Gold 6342 CPUs (2.80 GHz, 32 GB system memory) and NVIDIA A100 GPUs (80 GB GPU memory). Training ALIGNED took an average of 10-12 hours per run. This demonstrates the practical scalability of ALIGNED on genome-scale problems with comprehensive knowledge bases.
D.2 Reproducibility and Hyper-parameters
Unless specified, our experiments on ALIGNED and other methods used random seeds to split for training, validation and test set.
In Section 4.1, key hyper-parameters of our ALIGNED method as follows:
$k=4$ in Equation 2.2.
${\bm{w}}=0.5\big(\sum\limits_{{\bm{x}},{\bm{y}}\in\langle{\bm{X}},{\bm{Y}}\rangle}\bm{1}_{{\bm{y}}=\delta_{\mathcal{KB}}({\bm{x}})}/|{\bm{X}}|\big)+0.2\big(\sum\limits_{1\leq i\leq n}({R}^{(0)}_{+,i,j}+{R}^{(0)}_{-,i,j})/2n\big)+0.3{\bm{c}}_{GO}$ in Equation 4, tuned on Norman et al. (2019) dataset and Türei et al. (2016) knowledge base, where $\langle{\bm{X}},{\bm{Y}}\rangle$ is the training dataset, $\mathcal{KB}=\langle{\bm{R}}_{+}^{(k)},{\bm{R}}_{-}^{(k)}\rangle$ is the knowledge base, $n$ is the total number of genes. ${\bm{c}}_{GO}$ is the number of GO annotations for each gene, normalized to $[0,1]$.
$\gamma=5$ in Equation 3; $C_{w}=10$, $C_{l}=5$, $\theta=0.3|{\bm{a}}|$ in Equation 4; $C=10$ in Equation 5; $\lambda=1$, $t_{0}=100$ and $t_{k}=1$ in Equation 6.
D.3 Full Results of Consistency Benchmark on Norman et al. (2019); Dixit et al. (2016); Adamson et al. (2016) Datasets
Table 2: Consistency benchmark in Figure 3,4
Dataset Model ABL Stage Data Cons. Knowledge Cons. Balanced Cons.
Norman linear - $0.4875_{\pm 0.0000}$ $0.3009_{\pm 0.0000}$ $0.3621_{\pm 0.0000}$
GEARS - $0.4755_{\pm 0.0063}$ $0.3175_{\pm 0.0100}$ $0.3731_{\pm 0.0066}$
scGPT - $0.4846_{\pm 0.0000}$ $0.3016_{\pm 0.0000}$ $0.3622_{\pm 0.0000}$
scFoundation - $0.4805_{\pm 0.0007}$ $0.3110_{\pm 0.0008}$ $0.3692_{\pm 0.0006}$
ALIGNED
baseline $0.5360_{\pm 0.0019}$ $0.3767_{\pm 0.0063}$ $0.4192_{\pm 0.0062}$
Align 1 $0.5040_{\pm 0.0106}$ $0.4307_{\pm 0.0197}$ $0.4575_{\pm 0.0125}$
Refine 1 $0.5252_{\pm 0.0042}$ $0.6549_{\pm 0.0044}$ $0.5697_{\pm 0.0024}$
Align 2 $0.5246_{\pm 0.0047}$ $0.6528_{\pm 0.0064}$ $0.5687_{\pm 0.0024}$
Refine 2 $0.5248_{\pm 0.0046}$ $0.6573_{\pm 0.0045}$ $0.5698_{\pm 0.0028}$
ALIGNED
baseline $0.5386_{\pm 0.0017}$ $0.3820_{\pm 0.0109}$ $0.4242_{\pm 0.0102}$
Align 1 $0.5154_{\pm 0.0163}$ $0.4133_{\pm 0.0366}$ $0.4447_{\pm 0.0260}$
Refine 1 $0.5270_{\pm 0.0023}$ $0.6555_{\pm 0.0021}$ $0.5714_{\pm 0.0022}$
Align 2 $0.5261_{\pm 0.0032}$ $0.6559_{\pm 0.0024}$ $0.5711_{\pm 0.0025}$
Refine 2 $0.5261_{\pm 0.0033}$ $0.6596_{\pm 0.0019}$ $0.5717_{\pm 0.0029}$
Dixit linear - $0.2827_{\pm 0.0000}$ $0.1432_{\pm 0.0000}$ $0.1807_{\pm 0.0000}$
GEARS - $0.4330_{\pm 0.0017}$ $0.2704_{\pm 0.0029}$ $0.3243_{\pm 0.0029}$
scGPT - $0.4361_{\pm 0.0001}$ $0.2804_{\pm 0.0001}$ $0.3335_{\pm 0.0001}$
scFoundation - $0.4335_{\pm 0.0004}$ $0.2722_{\pm 0.0010}$ $0.3260_{\pm 0.0009}$
ALIGNED
baseline $0.4207_{\pm 0.0028}$ $0.3930_{\pm 0.0184}$ $0.4052_{\pm 0.0113}$
Align 1 $0.4126_{\pm 0.0022}$ $0.4245_{\pm 0.0168}$ $0.4173_{\pm 0.0069}$
Refine 1 $0.4188_{\pm 0.0039}$ $0.6591_{\pm 0.0036}$ $0.4718_{\pm 0.0041}$
Align 2 $0.4181_{\pm 0.0036}$ $0.6579_{\pm 0.0042}$ $0.4711_{\pm 0.0038}$
Refine 2 $0.4180_{\pm 0.0036}$ $0.6630_{\pm 0.0014}$ $0.4711_{\pm 0.0037}$
ALIGNED
baseline $0.4509_{\pm 0.0068}$ $0.3808_{\pm 0.0033}$ $0.4069_{\pm 0.0040}$
Align 1 $0.4534_{\pm 0.0102}$ $0.4147_{\pm 0.0251}$ $0.4289_{\pm 0.0126}$
Refine 1 $0.4585_{\pm 0.0048}$ $0.6367_{\pm 0.0122}$ $0.5082_{\pm 0.0046}$
Align 2 $0.4588_{\pm 0.0045}$ $0.6376_{\pm 0.0130}$ $0.5082_{\pm 0.0046}$
Refine 2 $0.4589_{\pm 0.0047}$ $0.6555_{\pm 0.0023}$ $0.5110_{\pm 0.0047}$
Adamson linear - $0.2806_{\pm 0.0000}$ $0.1866_{\pm 0.0000}$ $0.2198_{\pm 0.0000}$
GEARS - $0.4687_{\pm 0.0018}$ $0.3737_{\pm 0.0045}$ $0.4132_{\pm 0.0024}$
scGPT - $0.4703_{\pm 0.0000}$ $0.3563_{\pm 0.0001}$ $0.4016_{\pm 0.0001}$
scFoundation - $0.4704_{\pm 0.0002}$ $0.3678_{\pm 0.0013}$ $0.4097_{\pm 0.0009}$
ALIGNED
baseline $0.2888_{\pm 0.0030}$ $0.3554_{\pm 0.0293}$ $0.3128_{\pm 0.0077}$
Align 1 $0.4767_{\pm 0.0108}$ $0.3725_{\pm 0.0315}$ $0.4063_{\pm 0.0281}$
Refine 1 $0.4682_{\pm 0.0102}$ $0.4111_{\pm 0.0261}$ $0.4325_{\pm 0.0179}$
Align 2 $0.4672_{\pm 0.0086}$ $0.6467_{\pm 0.0056}$ $0.5178_{\pm 0.0078}$
Refine 2 $0.4674_{\pm 0.0078}$ $0.6418_{\pm 0.0083}$ $0.5178_{\pm 0.0077}$
ALIGNED
baseline $0.5436_{\pm 0.0059}$ $0.4319_{\pm 0.0317}$ $0.4692_{\pm 0.0268}$
Align 1 $0.4925_{\pm 0.0450}$ $0.5589_{\pm 0.1079}$ $0.5023_{\pm 0.0226}$
Refine 1 $0.5225_{\pm 0.0177}$ $0.6321_{\pm 0.0097}$ $0.5641_{\pm 0.0142}$
Align 2 $0.5205_{\pm 0.0168}$ $0.6328_{\pm 0.0105}$ $0.5611_{\pm 0.0127}$
Refine 2 $0.5235_{\pm 0.0179}$ $04_{\pm 0.0026} 0.5665±0$$65_{\pm 0.0143}$